# Automated Machine Learning Pipeline for Training and Analysis Using Large Language Models

### v1

@@ Abstract
Machine learning interatomic potentials (MLIPs) have become powerful tools to extend molecular simulations beyond the limits of quantum methods, offering near-quantum accuracy at much lower computational cost. Yet, developing reliable MLIPs remains difficult because it requires generating high-quality datasets, preprocessing atomic structures, and carefully training and validating models.
In this work, we introduce an Automated Machine Learning Pipeline (AMLP) that unifies the entire workflow from dataset creation to model validation. AMLP employs large-language-model agents to assist with electronic-structure code selection, input preparation, and output conversion, while its analysis suite (AMLP-Analysis) based on ASE supports a range of molecular simulations. The pipeline is built on the MACE architecture and validated on acridine polymorphs, where with a straightforward fine-tuning of a foundation model mean absolute errors of  1.7 meV/atom in energies and  7.0 meV/Å in forces are achieved. The fitted MLIP reproduces DFT geometries with sub-Å accuracy and demonstrates stability during molecular dynamics simulations in the microcanonical and canonical ensemble.
@@ 1 Introduction
Molecular dynamics (MD) simulations are primarily employed to investigate finite-temperature effects, dynamical properties, and time-correlation functions in materials.[1, 2] While classical force fields enable access to long time and length scales, their inherent approximations can limit the accuracy of these predictions.[3] These classical force fields are often parameterized or fitted to specific types of systems or interactions, leading to inadequacies when attempting to generalize their use to broader, more complex scenarios. Particularly problematic is their inability to reliably distinguish subtle energy differences between closely related polymorphs, which typically differ by only a few kJ/mol.[4, 5] Such precision is critical in determining material stability, phase transitions, and reactive pathways, highlighting the need for methods that can accurately describe these delicate energy landscapes.
@@
Quantum mechanical (QM) methods, such as Density Functional Theory (DFT) and post-Hartree-Fock approaches, offer high accuracy in capturing electronic structure details. However, their computational expense rapidly escalates with system size, making them impractical for routine simulations of large-scale molecular systems simulations.[6] This computational constraint limits their applicability to small or medium-sized systems.
@@
In response to these limitations, machine learning interatomic potentials (MLIPs) have emerged as a robust alternative, bridging the accuracy of quantum chemical methods and the computational efficiency of classical force field potentials.[7, 8] MLIPs learn from quantum chemical data, providing high-fidelity representations of atomic interactions while remaining computationally efficient enough to handle large systems and dynamical simulations over longer timescales. They offer substantial improvements in both predictive accuracy and scalability, facilitating precise simulations of complex systems under a wide variety of conditions. However, having these benefits depends on the quality and comprehensiveness of the training dataset.
@@
The key factor in developing an accurate model capable of effectively predicting user requirements lies in constructing a precise dataset. Initially, careful consideration of the underlying chemistry of the system is essential, particularly identifying the dominant interactions influencing system stability, such as long-range interactions for example. Subsequently, selecting the appropriate QM method is crucial. For instance, employing a plane-wave basis set is preferable over Gaussian basis sets when investigating crystalline structures, due to its better suitability for periodic systems.
@@
Furthermore, it is crucial to ensure that the dataset contains structural configurations extending beyond the equilibrium minima. Typically, this sampling procedure begins with geometry optimization (or cell optimization in the case of crystals systems), followed by ab initio molecular dynamics (AIMD) simulations conducted at various temperatures. In principle, such a protocol enables the dataset to sample a broad region of the potential energy surface (PES), capturing not only near-equilibrium configurations but also thermally accessible distortions, anharmonic vibrations, and local structural rearrangements. In practice, however, the extent of PES coverage is constrained by the timescales and temperature ranges explored during AIMD. For example, simulations at moderate temperatures primarily sample configurations near the equilibrium minima, whereas high-temperature trajectories can probe a wider basin of the PES, although at the cost of introducing highly distorted structures that may not be representative of typical conditions. The inclusion of these diverse structural samples significantly enhances the MLIP’s ability to generalize and adapt effectively across different simulation regimes, especially in MD scenarios. Conversely, a limited or inadequately constructed dataset covers only a narrow region of the PES, which can lead the model to learn spurious correlations and produce unrealistic or unreliable predictions outside the training domain.
Once the dataset is sufficiently comprehensive to handle diverse conditions, attention can be turned to selecting appropriate parameters for the training procedure. Recent developments in machine learning potentials increasingly rely on Graph Neural Networks (GNNs), which provide a natural way to represent atomic environments and have shown clear improvements in accuracy compared to earlier approaches.[9] In particular, the use of equivariant architectures—ensuring predictions respect rotational and translational symmetries—has further enhanced the reliability of these models.[10] To systematically evaluate such methods, large-scale benchmarking initiatives have emerged. The Matbench test suite provided a standardized benchmarks for inorganic materials properties,[11] and more recently, bench Discovery introduced a framework to assess machine learning predictions of crystal stability across vast chemical and structural spaces.[12] Extensive research efforts have already been devoted to developing large-scale benchmarks for molecular crystals, exemplified by the OMC25 dataset and its associated models, which aim to reproduce molecular properties with high fidelity.[13]
Examples of GNN-based MLIPs, such as MACE, have gained considerable attention.[14] MACE distinguishes itself through detailed modeling of atomic environments and provides foundational models pretrained across a broad spectrum of chemical systems. This approach significantly enhances predictive accuracy and reduces computational resources required during fine-tuning to specific use cases.[14, 15]
@@
To address these challenges, we introduce a comprehensive pipeline called Automated Machine Learning Pipeline (AMLP). AMLP simplifies the creation of MLIPs from straightforward input files, such as Crystallographic Information Files (.cif) or Cartesian coordinate files (.xyz). The AMLP pipeline comprises five distinct steps, each designed to streamline the overall process. Moreover, AMLP leverages Large Language Models (LLMs) to assist users in identifying suitable QM methodologies for their systems. In practice, the user provides a concise prompt describing the target system, and the LLM queries its knowledge of the published literature to suggest methods, basis sets, or dispersion corrections that have been successfully applied to similar cases. This capability not only offers an informed starting point for parameter selection but also helps the user rapidly survey prior research, thereby reducing the need for extensive manual literature review and enabling faster setup of reliable QM calculations. Subsequently, AMLP facilitates the automated creation of input files necessary for geometry optimization and AIMD simulations compatible with popular computational software such as Gaussian [16], VASP [17, 18, 19, 20], and CP2K.[21]
@@
Furthermore, the AMLP methodology is designed to be user-friendly and accessible, requiring minimal coding expertise. Users simply specify a numerical input corresponding to the desired stage of the workflow, ensuring a straightforward and intuitive experience. The resulting potentials can subsequently be utilized through an analysis module, AMLP-Analysis, integrated with the Python-based Atomic Simulation Environment (ASE), facilitating flexible evaluation and simulation tasks.[22]
@@ 2 Code architecture
2.1 General flow chart of the AMLP pipeline
@@
A tutorial that guides users through the general workflow is available online.[23] Even non-expert users can follow the steps to process single .cif files and construct a machine learning potential within a single session.
2.1.1 AI-agent feedback
The training procedure begins with AI-driven agents designed to provide users with tailored insights into their chemical systems, in line with their specific research objectives. These agents also offer an initial suggestion regarding the appropriate DFT code and parameter settings that may be suitable for their study. While LLMs are already being explored in the field of chemistry to generate insights from user queries, this area is still rapidly evolving due to its precision and ease of integration in different codes and workflows.[24]
Numerous efforts are underway to integrate LLMs into domain-specific computational frameworks.[25, 26] In our approach, we follow a similar paradigm by using an LLM-based model to suggest initial DFT setups. Additionally, we introduce an application of general-purpose AI tools, where queries are progressively refined and delegated to specialized agents tailored.
@@ 2.1.1 AI-agent feedback
The literature review is implemented through the PublicationAPI class, which retrieves and analyzes publications from multiple databases, including arXiv[27], Crossref[28], PubMed[29], and Semantic Scholar[30]. These sources can be modified in the same configuration file where the agents’ models are defined. Parameter extraction relies primarily on LLM-based methods, with a fallback to regular expressions (Regex) if LLM parsing fails. Regex provides pattern-matching tools that locate, extract, and validate text based on predefined syntax. In this context, Regex serves as a backup for extracting computational details—such as DFT parameters—from scientific literature when AI-based extraction is unsuccessful.
@@ 2.1.4 AIMD processing
Although AIMD simulations are powerful tools for sampling diverse molecular configurations, it is strongly recommended to perform a post-simulation validation of selected structures using quantum chemical methods (e.g., Gaussian or VASP single-point calculations). This step ensures that the energies and geometries obtained from the AIMD trajectory are physically meaningful and suitable for further analysis or training purposes.
@@
2.1.5 ML potential dataset creation
@@
The dataset containing all relevant information in a single .json file can be automatically converted into an HDF5 format, which is then partitioned into training and validation subsets—by default using an 85% to 15% split, although this ratio can be adjusted by the user. During this process, users are also prompted to define the batch size and specify a maximum threshold for atomic forces, which helps tailor the dataset to the intended training conditions. Any configurations exceeding this force threshold are automatically identified and excluded as outliers, thereby improving the robustness and physical relevance of the final dataset. Once the dataset is prepared, users can directly launch the training of a MLIP using the MACE architecture. A sample config.yaml file—containing all necessary parameters for training—is available in the project’s GitHub repository and can be adapted by users to suit their specific needs.[31] This ensures a straightforward transition from dataset generation to model training while preserving flexibility for methodological adjustments. Upon successful training, users obtain a customized MLIP, which can subsequently be directly loaded and utilized with the next module.
2.1.6 AMLP-A - Validation and Testing
@@
This subsection enables users to load and apply their MLIP for a variety of computational analyses. The process is streamlined through a straightforward config.yaml configuration file, allowing users to conveniently execute multiple computational tasks simultaneously. This validation module uses ASE, a widely-used Python-based toolkit known for its simplicity, flexibility, and extensive range of features suitable for atomic-scale simulations.[22]
@@
Users have several options for simulation types, including single-point energy calculations, geometry optimizations and cell optimizations. Geometry optimization can be activated by setting the corresponding flag to in the config.yaml file. A variety of optimization algorithms provided by ASE, such as BFGS[32, 33, 34, 35], LBFGS[36], and FIRE[37], are available for selection based on user preferences and requirements. Detailed information about these optimizers can be readily accessed through the ASE documentation.[38] Users must define an appropriate convergence criterion ($f_{max}$) to guide the optimization process. Upon successful completion of geometry optimization, the resulting data—including lattice parameters, total energy, simulation cell dimensions and angles, stress tensors, atomic coordinates, and individual atomic forces—are comprehensively reported and systematically saved into an .xyz file for further computational tasks.
@@
Subsequent MD simulations can be performed. For simulations in the canonical (NVT) ensemble, users may employ either a Langevin[39] or a Nosé-Hoover thermostat[40, 41]. The Langevin integrator is particularly robust in dissipating excess energy and equilibrating systems with rugged potential energy surfaces, while the Nosé-Hoover chain integrator provides smoother canonical sampling without stochastic noise.
@@
Alternatively, simulations may be conducted in the microcanonical (NVE) ensemble, in which the total energy is conserved. Here, the velocity-Verlet integrator is used to propagate the dynamics. Users can therefore choose between deterministic (Nosé-Hoover, velocity-Verlet) and stochastic (Langevin) schemes depending on the desired ensemble and level of control over thermal fluctuations. Energy conservation metrics are computed automatically for NVE runs to verify the stability of the integration. The energy conservation measure is calculated via
@@
where $E(0)$ is the reference energy at the beginning of the analysis period (after equilibration), $E_{k}$ is the total energy at step $k$, and $N_{step}$ is the number of simulation steps up to the current point. This formulation provides a cumulative measure of energy conservation quality that accounts for all deviations from the initial energy, normalized by the reference energy.[42] Additionally, users can replicate the computational cell in multiple spatial dimensions to model larger systems and achieve more comprehensive statistical sampling.
From the final MD trajectories, the radial distribution function (RDF) $g(r)$ is computed using the mdtraj library.[43] The RDF is defined as

### v2

@@ Abstract
Machine learning interatomic potentials (MLIPs) have become powerful tools to extend molecular simulations beyond the limits of quantum methods, offering near-quantum accuracy at much lower computational cost. Yet, developing reliable MLIPs remains difficult because it requires generating high-quality datasets, preprocessing atomic structures, and carefully training and validating models. In this work, we introduce an Automated Machine Learning Pipeline (AMLP) that unifies the entire workflow from dataset creation to model validation. AMLP employs large-language-model agents to assist with electronic-structure code selection, input preparation, and output conversion, while its analysis suite (AMLP-Analysis) based on ASE supports a range of molecular simulations. The pipeline is built on the MACE architecture and validated on acridine polymorphs, where with a straightforward fine-tuning of a foundation model mean absolute errors of  1.7 meV/atom in energies and  7.0 meV/Å in forces are achieved. The fitted MLIP reproduces DFT geometries with sub-Å accuracy and demonstrates stability during molecular dynamics simulations in the microcanonical and canonical ensemble.
@@ 1 Introduction
Molecular dynamics (MD) simulations are primarily employed to investigate finite-temperature effects, dynamical properties, and time-correlation functions in materials.[Md-headtransfer, MD-temperature-elastic] While classical force fields enable access to long time and length scales, their inherent approximations can limit the accuracy of these predictions.[limitation-MD] These classical force fields are often parameterized or fitted to specific types of systems or interactions, leading to inadequacies when attempting to generalize their use to broader, more complex scenarios. Particularly problematic is their inability to reliably distinguish subtle energy differences between closely related polymorphs, which typically differ by only a few kJ/mol.[md-poly1, md-poly2] Such precision is critical in determining material stability, phase transitions, and reactive pathways, highlighting the need for methods that can accurately describe these delicate energy landscapes.
@@
Quantum mechanical (QM) methods, such as Density Functional Theory (DFT) and post-Hartree-Fock approaches, offer high accuracy in capturing electronic structure details.[qc-expense] However, their computational expense rapidly escalates with system size, making them impractical for routine simulations of large-scale molecular systems simulations. This computational constraint limits their applicability to small or medium-sized systems.
@@
In response to these limitations, machine learning interatomic potentials (MLIPs) have emerged as a robust alternative, bridging the accuracy of quantum chemical methods and the computational efficiency of classical force field potentials.[QC-ML1, QC-ML2] MLIPs learn from quantum chemical data, providing high-fidelity representations of atomic interactions while remaining computationally efficient enough to handle large systems and dynamical simulations over longer timescales. They offer substantial improvements in both predictive accuracy and scalability, facilitating precise simulations of complex systems under a wide variety of conditions. However, having these benefits depends on the quality and comprehensiveness of the training dataset.
@@
The key factor in developing an accurate model capable of effectively predicting user requirements lies in constructing a precise dataset. Initially, careful consideration of the underlying chemistry of the system is essential, particularly identifying the dominant interactions influencing system stability, such as long-range interactions for example. Subsequently, selecting the appropriate QM method is crucial. While DFT remains the most common training data source for MLIPs due to its efficiency and moderate computational cost, higher-level QM methods, including CCSD [ccsdpaper-reviewer1] and multi-reference approaches,[multireference-reviewer1] can be used as the target level of theory.
@@
Furthermore, it is crucial to ensure that the dataset contains structural configurations extending beyond the equilibrium minima. Typically, this sampling procedure begins with geometry optimization (or cell optimization in the case of crystals systems), followed by ab initio molecular dynamics (AIMD) simulations conducted at various temperatures. In principle, such a protocol enables the dataset to sample a broad region of the potential energy surface (PES), capturing not only near-equilibrium configurations but also thermally accessible distortions, anharmonic vibrations, and local structural rearrangements. In practice, however, the extent of PES coverage is constrained by the timescales and temperature ranges explored during AIMD. For example, simulations at moderate temperatures primarily sample configurations near the equilibrium minima, whereas high-temperature trajectories can probe a wider basin of the PES, although at the cost of introducing highly distorted structures that may not be representative of typical conditions. The inclusion of these diverse structural samples significantly enhances the MLIP’s ability to generalize and adapt effectively across different simulation regimes, especially in MD scenarios. Conversely, a limited or inadequately constructed dataset covers only a narrow region of the PES, which can lead the model to learn spurious correlations and produce unrealistic or unreliable predictions outside the training domain. Standardized datasets now support evaluation across distinct domains—for example, Matbench for inorganic materials, Matbench Discovery for large-scale crystal-stability assessment, and OMC25 for molecular crystals.[mathbench, MatbenchDiscovery, omc25] However, for system-specific training, these datasets must be expanded. To address this, our Automated Machine Learning Pipeline (AMLP) is designed to function as a “data engine” within a human-driven active-learning (AL) framework for MLIPs. This design follows established users total control over each step of the AL cycle. The result is a portable, adjustable workflow—transferable across new structures and DFT codes, from data selection through final validation. To make this workflow accessible to non-experts, we integrated an LLM-powered multi-agent system that provides guidance on system-specific problems. This approach, which is under intense development in the community, has shown promise in solving direct quantum-chemistry tasks and facilitating ad hoc problem solving.[ZOU2025102263] Once a dataset is assembled, model selection and hyperparameter optimization become central to reliable training and fair comparison. For the training stage, MACE is a strong choice: it offers pretrained foundation models spanning broad chemical space with a design that captures fine-grained atomic environments, delivering strong baseline accuracy while markedly reducing data and compute requirements during task-specific fine-tuning [MACE, fine-tuning-juttafriend].
@@
To address these challenges, we introduce a comprehensive pipeline called Automated Machine Learning Pipeline (AMLP). AMLP simplifies the creation of MLIPs from straightforward input files, such as Crystallographic Information Files (.cif) or Cartesian coordinate files (.xyz). AMLP comprises five distinct steps, each designed to streamline the overall process. Moreover, AMLP leverages Large Language Models (LLMs) to assist users in identifying suitable QM methodologies for their systems. In practice, the user provides a concise prompt describing the target system, and the LLM queries its knowledge of the published literature to suggest methods, basis sets, or dispersion corrections that have been successfully applied to similar cases. This capability not only offers an informed starting point for parameter selection but also helps the user rapidly survey prior research, thereby reducing the need for extensive manual literature review and enabling faster setup of reliable QM calculations. Subsequently, AMLP facilitates the automated creation of input files necessary for geometry optimization and AIMD simulations compatible with popular computational software such as Gaussian[g16], VASP[vasp1, vasp2, vasp3, vasp4], CP2K[cp2K], Orca[orca], and Psi4[psi4].
@@
The resulting potentials can subsequently be utilized through an analysis module, AMLP-Analysis, integrated with the Python-based Atomic Simulation Environment (ASE), facilitating flexible evaluation and simulation tasks.[ase-paper] AMLP supports several dataset-construction strategies to achieve comprehensive coverage of the PES and minimize unsampled regions. Users may (i) perform cell optimizations followed by AIMD to generate training configurations and train an in-house MLIP; (ii) leverage foundation models to explore configuration space via the AMLP-Analysis module, then compute single-point reference data for the sampled structures using the data-generation module; or (iii) adopt a hybrid, active-learning workflow that begins with low-cost methods and iteratively refines the model with selectively added high-level calculations. In all cases, AMLP provides a unified, user-friendly interface that enables practitioners to adopt any of these strategies with minimal effort.
@@ 2 Code architecture
2.1 General flow chart of AMLP
@@
A tutorial that guides users through the general workflow is available online.[colablink-amlp] Even non-expert users can follow the steps to process single .cif files and construct a machine learning potential within a single session.
2.2 Part A: Data preparation pipeline
  1. 1.
    AI-agent feedback
    The training procedure begins with AI-driven agents designed to provide users with tailored insights into their chemical systems, in line with their specific research objectives. These agents also offer an initial suggestion regarding the appropriate DFT code and parameter settings that may be suitable for their study. While LLMs are already being explored in the field of chemistry to generate insights from user queries, this area is still rapidly evolving due to its precision and ease of integration in different codes and workflows.[comparison_model_studens]
    Numerous efforts are underway to integrate LLMs into domain-specific computational frameworks.[researchassi, chemgraph] In our approach, we follow a similar paradigm by using an LLM-based model to suggest initial DFT setups. Additionally, we introduce an application of general-purpose AI tools, where queries are progressively refined and delegated to specialized agents tailored.
@@ 2.2 Part A: Data preparation pipeline
    The literature review is implemented through the PublicationAPI class, which retrieves and analyzes publications from multiple databases, including arXiv[arxiv], Crossref[crossref], PubMed[pubmed], ChemRxiv [chemrxiv] and Semantic Scholar[semantic_scholar]. These sources can be modified in the same configuration file where the agents’ models are defined. Parameter extraction relies primarily on LLM-based methods, with a fallback to regular expressions (Regex) if LLM parsing fails. Regex provides pattern-matching tools that locate, extract, and validate text based on predefined syntax. In this context, Regex serves as a backup for extracting computational details—such as DFT parameters—from scientific literature when AI-based extraction is unsuccessful.
@@
  5. 5.
    ML potential dataset creation
@@
    The dataset containing all relevant information in a single .json file can be automatically converted into an HDF5 format, which is then partitioned into training and validation subsets—by default using an 85% to 15% split, although this ratio can be adjusted by the user. During this process, users are also prompted to define the batch size and specify a maximum threshold for atomic forces. While including moderately “high-force” configurations is beneficial for training robust MLIPs (as discussed in Section 1), excessively high-force configurations—often artifacts of insufficient equilibration, numerical instabilities, or unphysical/inaccessible geometries—can corrupt the learning of the PES and lead to poor generalization. The force threshold allows users to identify and exclude such extreme outliers while retaining the diverse, physically relevant configurations that improve model performance.” Once the dataset is prepared, users can directly launch the training of a MLIP using the MACE architecture. A sample config.yaml file—containing all necessary parameters for training—is available in the project’s GitHub repository and can be adapted by users to suit their specific needs.[amlp_github] This ensures a straightforward transition from dataset generation to model training while preserving flexibility for methodological adjustments.
The user then initiates MACE training using the AMLP-generated HDF5 datasets and configuration templates. Upon successful training, users obtain a customized MLIP, which can subsequently be directly loaded and utilized with the next module.
2.3 Part B: Post-training simulation tools
@@
This subsection enables users to load and apply their MLIP for a variety of computational analyses using the AMLP-Analyse (AMLPA) module. The process is streamlined through a straightforward config.yaml configuration file, allowing users to conveniently execute multiple computational tasks simultaneously. This validation module uses ASE, a widely-used Python-based toolkit known for its simplicity, flexibility, and extensive range of features suitable for atomic-scale simulations.[ase-paper]
@@
Users have several options for simulation types, including single-point energy calculations, geometry optimizations and cell optimizations. Geometry optimization can be activated by setting the corresponding flag to in the config.yaml file. Detailed information about these optimizers can be readily accessed through the ASE documentation.[ase_website] Upon successful completion of geometry optimization, the resulting data—including lattice parameters, total energy, simulation cell dimensions and angles, stress tensors, atomic coordinates, and individual atomic forces—are comprehensively reported and systematically saved into an .xyz file for further computational tasks.
@@
Subsequent MD simulations can be performed. For simulations in the canonical (NVT) ensemble, users may employ either a Langevin[SchneiderStoll1978] or a Nosé-Hoover chain thermostat[Nose1984, Hoover1985, Martyna1992]. The Langevin integrator is particularly robust in dissipating excess energy and equilibrating systems with rugged potential energy surfaces, while the Nosé-Hoover chain integrator provides smoother canonical sampling without stochastic noise.
@@
Alternatively, simulations may be conducted in the microcanonical (NVE) ensemble, in which the total energy is conserved. Here, the velocity-Verlet integrator is used to propagate the dynamics. Users can therefore choose between deterministic (Nosé-Hoover chain, velocity-Verlet) and stochastic (Langevin) schemes depending on the desired ensemble and level of control over thermal fluctuations. Energy conservation metrics are computed automatically for NVE runs to verify the stability of the integration. The energy conservation measure is calculated via
@@
where $E(0)$ is the reference energy at the beginning of the analysis period (after equilibration), $E_{k}$ is the total energy at step $k$, and $N_{step}$ is the number of simulation steps up to the current point. This formulation provides a cumulative measure of energy conservation quality that accounts for all deviations from the initial energy, normalized by the reference energy.[marksbook] Additionally, users can replicate the computational cell in multiple spatial dimensions to model larger systems and achieve more comprehensive statistical sampling.
From the final MD trajectories, the radial distribution function (RDF) $g(r)$ is computed using the mdtraj library.[mdtraj] The RDF is defined as